OpenStudy (anonymous):

Write the equation of the line that passes through (2, 2) and (6, 3) in standard form. 4x + 4y = 6 x + 4y = 12 4x + 4y = –12 x – 4y = –6

OpenStudy (anonymous):

@zepp

OpenStudy (anonymous):

@ zepp will tell you how to do this..

OpenStudy (zepp):

Find the slope!

OpenStudy (zepp):

\[\large m=\frac{y_2-y_1}{x_2-x_1}\] Using the two points, (2, 2) and (6, 3) \[\large m=\frac{3-2}{6-2}=\frac{1}{4}\]

OpenStudy (zepp):

Then use the point slope formula @waterineyes showed you earlier, now it's your turn :D

OpenStudy (zepp):

\[\large y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

i just tried it and it didnt give me any of the answers above!

OpenStudy (anonymous):

Show your work @priscillaam_0422 So that we can find your mistake..

OpenStudy (zepp):

Show us what you did and we'll try to find where you made a mistake.

OpenStudy (anonymous):

Good Work @zepp ..